Y is a random variable that is distributed N(10,324). Find Prob(Y < 13.42 OR Y > 34.3).
Answer #1

YONC10, 324) =) N=10, o = √324 = 18 P(Y213.42 or y334.3) = P(Y<13-42) + P (Y334-3) = P (Z<0-17) + P(Z31:35) = 0.5753 + 0.0885